# Integrated Open-Source Framework for Simulation of Transcatheter Pulmonary Valves in Native Right Ventricular Outflow Tracts

### v2

@@ 3.4.6 Simulation of Device Expansion
The simulation was divided into two analysis steps to capture the curved compression of the device, as though fitting it into the catheter delivery system, followed by the staged deployment in the vessel, as also shown in Figure 4. The distal half of the TPV device expands before the proximal half to mimic the delivery system sheath being pulled back and the device released. The tube was compressed over one second at a rate such that the enclosed device was well contained within the vessel. To achieve this, a displacement normal to the surface of the curved tube was applied to compress the device uniformly. Although the devices should fit within a delivery catheter system with a diameter of 8.33 mm (25 Fr) [29], we experienced instability in the models at very high compression due to complexities in the contact problem, so we used a larger target diameter of 16 mm after compression for this study. In most of our example cases, the compression rate was 20 mm/s, for the tube had an initial radius of 28 mm. When a greater initial radius was required for the tube to fully surround the device without penetration, the rate was adjusted accordingly to meet the same target radius. During TPV deployment, each half of the device was allowed to expand at the same rate at which it was compressed. The device compression and expansion simulation in FEBio is shown in Video-1.
@@
$E_{\alpha}=\kappa\left(\tilde{I}_{1}-3\right)+(1-3\kappa)\left(\tilde{I}_{4\alpha}-1\right)$ represents the fiber strain [37, 38]. For parameters, $k_{1}$ is the fiber modulus, $k_{2}$ is the fiber exponential coefficient, $k$ is the bulk modulus, $\kappa$ is the fiber dispersion, and $c$ is the shear modulus of the ground matrix. We used HGO material properties as defined in Donahue et al. [39].
We adopted the isotropic elastic material model [40] present in FEBio for the Nitinol TPV. The density, Young’s modulus $E$, and Poisson’s ratio $\nu$ must be defined. We chose known material properties for Nitinol [41]. The Nitinol device had a density of 6.45e-06 kg/mm3 and a Poisson’s ratio of 0.33. Nitinol exhibits superelasticity, which means that its stiffness transitions from a stiff austenite phase to a softer martensite phase to accommodate the large strains experienced during compression. When a stent expands after being unloaded, the material undergoes a reverse transformation from martensite back to austenite, resulting in an increase in stiffness as it returns to its original shape. As such, a step function load curve was applied to the Young’s modulus to simulate the transition of the effective stiffness in the device between the compression and release phases. The hyperelastic strain-energy function for the isotropic elastic material is given as:
@@
We used an unconstrained Neo-Hookean material [42] for the tube that compresses the device. For simplicity, the tube was assigned parameter values identical to the TPV, but the Young’s modulus was kept constant at 40 GPa. The strain-energy function is
@@
where $C$ is the right Cauchy-Green deformation tensor, $I_{1}$ and $I_{2}$ are the first and second invariants of $C$, and $J$ is the determinant of the deformation gradient tensor. The model captures the non-linear stress-strain response but reduces to classical linear elasticity under small strains and small rotations [43].
Boundary conditions and contacts were imposed on all geometries but varied by step. A potential-based contact formulation was used to constrain the TPV to stay within the tube during compression and staged expansion [44, 45]. The same contact formulation was used between the TPV and the vessel during deployment of the device within the RVOT. A zero displacement boundary condition was imposed on the central nodes of the TPV to prevent sliding along the vertical Z axis. The ends of the vessel were fixed. Expansion of the TPV was controlled by a normal displacement boundary condition applied to the surrounding tube and by the enforced contact potential, mimicking the self-expanding behavior of the Nitinol devices. To improve the stability of the simulation, mass damping was assigned to the device as part of a false transient approach. The vessel is pressurized in the diastolic CT images from which we made the mesh of the RVOT, so it needs to be pressurized in the simulation as well. Therefore, we applied a 10 mmHg pressure to the luminal surface of the vessel. Pulmonary artery diastolic pressure normally falls within the range of about 4 mmHg to 16 mmHg, so we chose the median value [46, 47]. The pressure ramped linearly over 1.0 s since the vessel was not in contact during the device compression phase of the simulation, which spans from 0 to 1.0 s. The pressure on the vessel wall then remained constant during deployment of the TPV.
@@ 4 RESULTS
Time taken for automatic segmentation of the RVOT models was 5 seconds. Time taken for device placement was 45 seconds. The Alterra model required more elements than the TPV-25 due to a more complex shape, which contributed to an elevated runtime.
@@
Across these three RVOT models, the 75th percentile strain and stress values were fairly similar between the two stents. For RVOT4, the TPV-25 resulted in greater 99th percentile and maximum strain and stress than the Alterra. For both stents, strain and stress at the proximal half of the RVOT were greater than those at the distal half. For RVOT12, like RVOT4, the TPV-25 produced greater 99th percentile and maximum strain and stress than the Alterra. However, distal strain was greater than the proximal strain. Proximal stress was greater than distal stress when the TPV-25 was deployed, while the reverse was true, i.e., distal stress was greater than proximal stress when the Alterra was deployed. In RVOT28, trends varied from RVOTs 4 and 12. At the proximal half, TPV-25 resulted in greater 99th percentile strain and stress, but lesser maximum strain and stress than the Alterra. At the distal half of the RVOT, TPV-25 produced a lesser 99th percentile and maximum strain and stress than the Alterra. On comparing the two halves, the distal half of RVOT28 showed greater strain and stress values than the proximal half across both stents, except for the maximum stress when the TPV-25 was used.
Contact maps generated for each RVOT and stent condition are shown in figure 7. Similar to the strain and stress distribution across the entire vessel, greater contact was observed at the proximal and distal ends of the RVOT where the stent geometries are the broadest. Overall, the TPV-25 had a greater percentage of the total stent area in contact with the vessel walls than the Alterra. On analyzing by location, the TPV-25 had fairly similar contact percentages in the proximal and distal halves. However, for RVOTs 4 and 12, the Alterra device contacted only  1% of its proximal area of the stent in the proximal half. For the Alterra stent, contact percentages in the distal half were much greater than those in the proximal half, unlike the TPV-25 stent.
By quantifying and comparing stress, strain, and contact patterns across multiple patient-specific anatomies and devices, this study underscores the importance of personalized planning in TPVR. The observed variability in these metrics across different RVOTs, between stents (TPV-25 vs. Alterra), and locations in the RVOT (proximal vs. distal) highlights that no single device performs optimally in all anatomies, reinforcing the need for simulation-driven preprocedural assessment.
@@ 5 DISCUSSION
While image-based simulations are increasingly used to support adult structural interventions, similar workflows have yet to be routinely adopted in congenital heart disease. [48, 49, 50] However, patient-specific modeling may be especially valuable in this setting due to the high anatomical variability and small patient cohorts, which limit the feasibility of large clinical trials and empirical design validation.
Several groups have explored the modeling of TPVR over the last two decades, which laid the groundwork for the current generation of patient-specific simulations. Capelli et al. [15, 51] pioneered early simulations of self-expanding devices in image-derived RVOT models, laying the foundation for later studies. Their early implementations assumed a rigid, non-deformable RVOT, now recognized as a significant simplification [10]. Bosi et al. [52] advanced the field by introducing deformable vessel behavior and simulating both balloon-expandable and self-expanding valves using Abaqus, though with simplified isotropic linear elastic models for the vessel wall. Further advancements have aimed to improve the physiological accuracy of vessel mechanics. Donahue et al. [39], for example, applied the Holzapfel-Gasser-Ogden (HGO) constitutive model to account for anisotropy and fiber orientation in RVOT tissue. However, their work focused on simulating balloon expansion through wall pressurization to assess coronary compression risk, rather than modeling the deployment of a self-expanding TPV device.
Notably, these foundational simulations predated the commercial availability of dedicated self-expanding TPV devices for pediatric patients. In this study, we address that gap by simulating the deployment of commercially available self-expanding TPVs in a range of pediatric RVOT anatomies. Using the HGO model to represent the RVOT wall, we demonstrate the feasibility of preprocedural prediction of device expansion, positioning, and vessel interaction in patient-specific cases. To further enhance the practicality of our workflow, we incorporated machine-learning-based segmentation for automated RVOT extraction. While now increasingly routine, this capability greatly improves efficiency and supports the clinical applicability of patient-specific modeling, particularly in complex congenital anatomies [27].
@@
Device–tissue interaction was modeled using a contact potential method, based on the formulation by Kamensky et al. [44], which applies a repulsive force when surfaces come within a specified distance. Although this approach maintains an artificial small gap between the stent and vessel wall, it prevents penetration and offers a more stable solution compared to the sliding-elastic formulation in FEBio, especially for soft tissue interactions [25].
Notably, the material phase transformation in the Nitinol stent is influenced by the progression of the martensitic volume fraction, which changes in response to internal stress during deformation. Although our step-function approach provides a reasonable approximation of the overall characteristics of the stent, it does not fully capture the mechanical hysteresis behavior associated with the transformation. Subsequent work will aim to implement a more physically representative superelasticity model [53] based on the martensite fraction, similar to those presented in previous literature [54, 55, 56]. A more representative material model for the stent may result in different softening and stiffening mechanisms that could influence the final expansion of the stent, and the strain and stress values and distribution in the RVOT. However, this needs to be explored further.
@@
In our simulations, the RVOT wall was modeled using HGO parameters from existing literature [39], without accounting for layer-specific material properties due to the lack of experimental data in pediatric tissues. Currently, tissue-derived material data for the pulmonary artery are limited to a small number of patients, ranging in age from infancy [57] to older adults [58]. Given the difficulty of in vivo mechanical characterization and expected interpatient variability of vessel wall parameters, simulation accuracy will continue to improve as new data and novel measurement techniques become available [59].
In this proof-of-concept study, we performed simulations only in the diastolic phase of the cardiac cycle. However, the RVOT and pulmonary artery are dynamic, and future simulations will include both diastolic and systolic models to allow comparison over the cardiac cycle.
Finally, another limitation of this study is the small sample size and lack of validation against clinical post-implantation geometries. Future work will include a larger cohort and comparisons with post-procedural imaging to evaluate predictive accuracy.
@@ 6 CONCLUSION
The open-source and adaptable nature of our platform makes it well-suited for continued advancement and broader application. By enabling the development and testing of patient-specific workflows, it provides a foundation for evaluating both novel devices and new uses of existing technologies [60]. With ongoing refinement and validation, this framework has the potential to become a valuable clinical tool for optimizing device selection and guiding procedural planning in TPVR.
7 FUNDING
This work was supported by NIH R01HL153166, K25HL168235, 2R01GM083925, a CHOP Cardiac Center Innovation Award, the Topolewski Pediatric Valve Center at CHOP, and a CHOP Research Institute Post-Frontier Award.
8 DISCLOSURES

### v3

@@ 3.4.6 Simulation of Device Expansion
The simulation was divided into two analysis steps to capture the curved compression of the device, as though fitting it into the catheter delivery system, followed by the staged deployment in the vessel, as also shown in Figure 4. The distal half of the TPV device expands before the proximal half to mimic the delivery system sheath being pulled back and the device released. The tube was compressed over one second at a rate such that the enclosed device was well contained within the vessel. To achieve this, a displacement normal to the surface of the curved tube was applied to compress the device uniformly. Although the devices should fit within a delivery catheter system with a diameter of 8.33 mm (25 Fr) [medtronic_tpv_2021], we experienced instability in the models at very high compression due to complexities in the contact problem, so we used a larger target diameter of 16 mm after compression for this study. In most of our example cases, the compression rate was 20 mm/s, for the tube had an initial radius of 28 mm. When a greater initial radius was required for the tube to fully surround the device without penetration, the rate was adjusted accordingly to meet the same target radius. During TPV deployment, each half of the device was allowed to expand at the same rate at which it was compressed. The device compression and expansion simulation in FEBio is shown in Video-1.
@@
$E_{\alpha}=\kappa\left(\tilde{I}_{1}-3\right)+(1-3\kappa)\left(\tilde{I}_{4\alpha}-1\right)$represents the fiber strain [gasser_hyperelastic_2006, maas_uncoupled_2025]. For parameters, $k_{1}$ is the fiber modulus, $k_{2}$ is the fiber exponential coefficient, $k$ is the bulk modulus, $\kappa$ is the fiber dispersion, and $c$ is the shear modulus of the ground matrix. We used HGO material properties as defined in Donahue et al. [donahue_finite_2022].
We adopted the isotropic elastic material model [maas_isotropic_2025] present in FEBio for the Nitinol TPV. The density, Young’s modulus $E$, and Poisson’s ratio $\nu$ must be defined. We chose known material properties for Nitinol [vemury_behaviour_2019]. The Nitinol device had a density of 6.45e-06 kg/mm3 and a Poisson’s ratio of 0.33. Nitinol exhibits superelasticity, which means that its stiffness transitions from a stiff austenite phase to a softer martensite phase to accommodate the large strains experienced during compression. When a stent expands after being unloaded, the material undergoes a reverse transformation from martensite back to austenite, resulting in an increase in stiffness as it returns to its original shape. As such, a step function load curve was applied to the Young’s modulus to simulate the transition of the effective stiffness in the device between the compression and release phases. The hyperelastic strain-energy function for the isotropic elastic material is given as:
@@
We used an unconstrained Neo-Hookean material [maas_neo-hookean_2025] for the tube that compresses the device. For simplicity, the tube was assigned parameter values identical to the TPV, but the Young’s modulus was kept constant at 40 GPa. The strain-energy function is
@@
where $C$ is the right Cauchy-Green deformation tensor, $I_{1}$ and $I_{2}$ are the first and second invariants of $C$, and $J$ is the determinant of the deformation gradient tensor. The model captures the non-linear stress-strain response but reduces to classical linear elasticity under small strains and small rotations [bonet_nonlinear_2008].
Boundary conditions and contacts were imposed on all geometries but varied by step. A potential-based contact formulation was used to constrain the TPV to stay within the tube during compression and staged expansion [kamensky_contact_2018, maas_contact_2025]. The same contact formulation was used between the TPV and the vessel during deployment of the device within the RVOT. A zero displacement boundary condition was imposed on the central nodes of the TPV to prevent sliding along the vertical Z axis. The ends of the vessel were fixed. Expansion of the TPV was controlled by a normal displacement boundary condition applied to the surrounding tube and by the enforced contact potential, mimicking the self-expanding behavior of the Nitinol devices. To improve the stability of the simulation, mass damping was assigned to the device as part of a false transient approach.
@@ 4 Results
Time taken for automatic segmentation of the RVOT models was 5 seconds. Time taken for device placement was 45 seconds. The average computational time was 2 hr 17 min for simulations with the TPV-25 and 5 hr 34 min for the Alterra on an Intel Core i9 10980 XE CPU @ 3.00GHz processor with 64.0 GB RAM system. The Alterra model required more elements than the TPV-25 due to a more complex shape, which contributed to an elevated runtime.
@@
Across these three RVOT models, the 75th percentile strain and stress values were fairly similar between the two stents. For RVOT4, the TPV-25 resulted in greater 99th percentile and maximum strain and stress than the Alterra. For both stents, strain and stress at the proximal and distal halves of the RVOT were similar. For RVOT12, like RVOT4, the TPV-25 produced a greater 99th percentile and maximum strain and stress than the Alterra. Proximal and distal strains and stresses were similar. In RVOT28, trends varied from RVOTs 4 and 12. At the proximal half, TPV-25 resulted in greater 99th percentile strain and stress, but lesser maximum strain and stress than the Alterra. At the distal half of the RVOT, TPV-25 produced a lower 99th percentile and maximum strain and stress than the Alterra. On comparing the two halves, the distal half of RVOT28 showed greater strain and stress values than the proximal half across both stents, except for the maximum stress when the TPV-25 was used.
Contact maps generated for each RVOT and stent condition are shown in figure 7. Greater contact was observed at the proximal and distal ends of the RVOT where the stent geometries are the broadest. Overall, the TPV-25 had a greater percentage of the total stent area in contact with the vessel walls than the Alterra. On analyzing by location, the TPV-25 had fairly similar contact percentages in the proximal and distal halves. However, for RVOTs 4 and 12, the Alterra device contacted only  1% and  6% of its proximal area of the stent in the proximal half. For the Alterra stent, contact percentages in the distal half were much greater than those in the proximal half, unlike the TPV-25 stent.
By quantifying and comparing stress, strain, and contact patterns across multiple patient-specific anatomies and devices, this study underscores the importance of personalized anatomical reconstruction in TPVR. The observed variability in these metrics across different RVOTs, between stents (TPV-25 vs. Alterra), and locations in the RVOT (proximal vs. distal) highlights that no single device performs optimally in all anatomies, reinforcing the need for simulation-driven preprocedural assessment.
@@ 5 Discussion
While image-based simulations are increasingly used to support adult structural interventions, similar workflows have yet to be routinely adopted in congenital heart disease. [xuan_stent_2017, brown_patient-specific_2023, grossi_validation_2024] However, patient-specific modeling may be especially valuable in this setting due to the high anatomical variability and small patient cohorts, which limit the feasibility of large clinical trials and empirical design validation.
Several groups have explored the modeling of TPVR over the last two decades, which laid the groundwork for the current generation of patient-specific simulations. Capelli et al. [capelli_patient-specific_2010, capelli_patient-specific_2012] pioneered early simulations of self-expanding devices in image-derived RVOT models, laying the foundation for later studies. Their early implementations assumed a rigid, non-deformable RVOT, now recognized as a significant simplification [jolley_toward_2019]. Bosi et al. [bosi_patientspecific_2015] advanced the field by introducing deformable vessel behavior and simulating both balloon-expandable and self-expanding valves using Abaqus, though with simplified isotropic linear elastic models for the vessel wall. Further advancements have aimed to improve the physiological accuracy of vessel mechanics. Donahue et al. [donahue_finite_2022], for example, applied the Holzapfel-Gasser-Ogden (HGO) constitutive model to account for anisotropy and fiber orientation in RVOT tissue. However, their work focused on simulating balloon expansion through wall pressurization to assess coronary compression risk, rather than modeling the deployment of a self-expanding TPV device.
Notably, these foundational simulations predated the commercial availability of dedicated self-expanding TPV devices for pediatric patients. In this study, we address that gap by simulating the deployment of commercially available self-expanding TPVs in a range of pediatric RVOT anatomies. Using the HGO model to represent the RVOT wall, we demonstrate the feasibility of preprocedural prediction of device expansion, positioning, and vessel interaction in patient-specific cases. To further enhance the practicality of our workflow, we incorporated machine-learning-based segmentation for automated RVOT extraction. While now increasingly routine, this capability greatly improves efficiency and supports the clinical applicability of patient-specific modeling, particularly in complex congenital anatomies [barak-corren_image-derived_2025].
@@
Device–tissue interaction was modeled using a contact potential method, based on the formulation by Kamensky et al. [kamensky_contact_2018], which applies a repulsive force when surfaces come within a specified distance. Although this approach maintains an artificial small gap between the stent and vessel wall, it prevents penetration and offers a more stable solution compared to the sliding-elastic formulation in FEBio, especially for soft tissue interactions [maas_febio_2012].
Notably, the material phase transformation in the Nitinol stent is influenced by the progression of the martensitic volume fraction, which changes in response to internal stress during deformation. Although our step-function approach provides a reasonable approximation of the overall characteristics of the stent, it does not fully capture the mechanical hysteresis behavior associated with the transformation. Subsequent work will aim to implement a more physically representative superelasticity model [auricchio_shape-memory_1997] based on the martensite fraction, similar to those presented in previous literature [auricchio_carotid_2011, tzamtzis_numerical_2013, maleckis_nitinol_2018]. A more representative material model for the stent may result in different softening and stiffening mechanisms that could influence the final expansion of the stent, and the strain and stress values and distribution in the RVOT. However, this needs to be explored further.
@@
In our simulations, the RVOT wall was modeled using HGO parameters from existing literature [donahue_finite_2022], without accounting for layer-specific material properties due to the lack of experimental data in pediatric tissues. Currently, tissue-derived material data for the pulmonary artery are limited to a small number of patients, ranging in age from infancy [cabrera_mechanical_2013] to older adults [holzapfel_determination_2005]. Non-conduit RVOTs, which include both native (non-patched) and surgically patched RVOTs, have anatomies that are categorized by substantial heterogeneity in geometry and biomechanical properties. These include unpredictable tissue compliance and distensibility. A comprehensive sensitivity analysis of RVOT material parameters may help understand the variation observed in RVOT behavior due to device expansion. Given the difficulty of in vivo mechanical characterization and expected interpatient variability of vessel wall parameters, simulation accuracy will continue to improve as new tissue mechanical testing data and novel measurement techniques become available [wu_noninvasive_2025].
CTA images, and hence, the RVOT geometries of all three patients, corresponded to the diastolic phase. Prior literature has reported that the phase of the CT dataset is crucial for assessing device stability [sturla_planning_2025]. While the use of the diastolic phase for analysis may have a negligible impact in the context of conduit RVOTs, its influence in non-conduit RVOTs, where compliance and dynamic dimensional changes are more pronounced, could be more substantial and needs to be explored further. Since the RVOT and pulmonary artery are dynamic, future simulations will include both diastolic and systolic models to allow comparison over the cardiac cycle.
Finally, another limitation of this study is the small sample size and lack of validation against clinical post-implantation geometries. Typically, post-implantation CTA scans in patients undergoing these procedures are performed only if there is a device functionality failure. Although the goal of this study was to present an end-to-end device compatibility workflow, future work will include a larger cohort and comparisons with post-procedural imaging to evaluate predictive accuracy.
@@ 6 Conclusion
The open-source and adaptable nature of our platform makes it well-suited for continued advancement and broader application. By enabling the development and testing of patient-specific workflows, it provides a foundation for evaluating both novel devices and new uses of existing technologies [tsao_transcatheter_2025]. With ongoing refinement and validation, this framework has the potential to become a valuable clinical tool for optimizing device selection and guiding procedural planning in TPVR.
[…]
Acknowledgements This work was supported by NIH R01HL153166, K25HL168235, 2R01GM083925, a CHOP Cardiac Center Innovation Award, the Topolewski Pediatric Valve Center at CHOP, and a CHOP Research Institute Post-Frontier Award.
Declarations
@@
Figure 1: Overview of Modeling Pipeline
Figure 2: A) Shell models of our 3 patient-specific RVOT vessels, to scale. The example vessels were chosen to demonstrate workflow applicability in a range of RVOT sizes and shapes. RVOT morphopolgy types are determined as per [schievano_variations_2007] B) TPV device models available in the TPVR Simulation module within SlicerHeart.
Figure 3: Device placement and TPVR simulation module functionality.
Figure 4: FEM modeling – staged deployment in FEBio.
Figure 5: A) 1st principal Lagrangian strain distribution in simulated RVOT and TPV cases, and B) peak strain values observed in the RVOT across simulated cases, separated by proximal and distal halves.
Figure 6: A) 1st principal stress distribution in simulated RVOT and TPV cases, and B) peak stress values observed in the RVOT across simulated cases, separated by proximal and distal halves.
Figure 7: Contact visualization for simulated RVOT and TPV cases, A) Visualization of the contact area on RVOT wall across all simulated cases, B) Bar plot depicting the percentage of the stent/valve frame in contact with the vessel wall, separated by proximal and distal halves.